OpenStudy (unicwaan):

Write the equation of the function g(x) if g(x) = f(x + 2) + 4 and f(x) = x3 – 7. May someone help me with this question for calculus? I don't quite understand it.

OpenStudy (unicwaan):

@SolomonZelman

OpenStudy (anonymous):

That means \[f(x)=x^3-7\] \[f(x+2)=(x+2)^3-7\] \[g(x)=(x+2)^{3}-7+4\]

OpenStudy (unicwaan):

Oh okay, that makes a lot of sense. Thank you!! @Shalante

OpenStudy (anonymous):

Glad you understood!